if the unit of force is 2N that of length and velocity is 6m/s then the unit mass is

Grade:11

4 Answers

Nicho priyatham
8 years ago
wt is the unit of length.......?
it is nt mentioned
Sunny Suri
8 years ago
sorry   length is 4 m
 
noogler
8 years ago
the unit of m is 2/9kg …........use F=mv2/r
Koolman
8 years ago
the unit of m is 2/9kg …........use F=mv2/r